A pulsed laser has an average output power of 12W. Each pulse consists of light at wavelength 400nm and has 16ms duration. Compute the number of photons that are emitted in a single pulse?
1
Expert's answer
2020-06-10T18:25:08-0400

The energy of laser beam

"E=P\tau=N\frac{hc}{\lambda}"

Hence, the number of photons in a single pulse


"N=\frac{\lambda P\tau}{hc}\\
=\frac{400\times 10^{-9}\times12\times 16\times 10^{-3}}{6.63\times 10^{-34}\times 3\times 10^8}=3.86\times 10^{17}"